# AttentionX: Exploiting Consensus Discrepancy In Attention from A Distributed Optimization Perspective

### v2

@@ 1 Introduction
In recent years, transformers [24] in artificial intelligence have achieved great success in various data-analysis domain such as natural language processing (NLP) [3; 23], computer vision [7], image generation and editing [20; 8; 29], and audio processing [13]. One key component in transformer is the Attention layer for capturing long-distance dependency across a sequence of tokens. Specifically, the Attention operation gathers relevant information from other tokens for each one via the $KQ$-softmax-based weighted summation of $V$-representations. The feedforward network (FFN) further processes the output of the Attention layer per token, which can be interpreted as local information fusion.
@@ 2 Brief Review of PDMM
Detailed convergence results of the algorithm can be found in [27; 21].

### v3

@@ 1 Introduction
In recent years, transformers [25] in artificial intelligence have achieved great success in various data-analysis domain such as natural language processing (NLP) [3; 24], computer vision [7], image generation and editing [21; 8; 30], and audio processing [13]. One key component in transformer is the Attention layer for capturing long-distance dependency across a sequence of tokens. Specifically, the Attention operation gathers relevant information from other tokens for each one via the $KQ$-softmax-based weighted summation of $V$-representations. The feedforward network (FFN) further processes the output of the Attention layer per token, which can be interpreted as local information fusion.
@@
We note that a recent work [20] (which was made public after our work) proposes to calculate the attention scores as the difference between two separate softmax attention maps before multiplying the $V$-representations. The purpose for doing so is to amplify attention to the relevant context while canceling noise. Even though our work and [20] have different motivations, from a high-level point of view, both approaches attempt to calculate the difference of two separate weighted summations of $V$-representations.
@@ 2 Brief Review of PDMM
Detailed convergence results of the algorithm can be found in [28; 22].